# The Curse of Depth in Large Language Models

### v1

@@ 6 Related Work
Layer Normalization in Language Models. LN (Ba, 2016) was initially applied after the residual connection in the original Transformer (Vaswani, 2017), which is known as Post-LN. Later on, Pre-LN (Baevski and Auli, 2019; Dai et al., 2019; Nguyen and Salazar, 2019) dominated LLMs, due to its compelling performance and stability (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023; Bi et al., 2024). Prior works have studied the effect of Pre-LN and Post-LN. Xiong et al. (2020) proves that Post-LN tends to have larger gradients near the output layer, which necessitates smaller learning rates to stabilize training, whereas Pre-LN scales down gradients with the depth of the model, working better for deep Transformers. Wang et al. (2019) empirically confirmed that Pre-LN facilitates stacking more layers and Post-LN suffers from gradient vanishing. The idea of connecting multiple layers was proposed in previous works (Bapna et al., 2018; Dou et al., 2018; Wang et al., 2019). Adaptive Model Initialization (Admin) was introduced to use additional parameters to control residual dependencies, stabilizing Post-LN. DeepNorm (Wang et al., 2024) enables stacking 1000-layer Transformer by upscaling the residual connection before applying LN. Additionally, Ding et al. (2021) proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. Takase et al. (2023a) introduced B2T to bypass all LN except the final one in each layer. Li et al. (2024b) recently combines Post-LN and Pre-LN to enhance the middle layers.
@@ 7 Conclusion
In this paper, we introduce the concept of the Curse of Depth in LLMs, highlighting an urgent yet often overlooked phenomenon: nearly half of the deep layers in modern LLMs are less effective than expected. We discover the root cause of this phenomenon is Pre-LN which is widely used in almost all modern LLMs. To tackle this issue, we introduce LayerNorm Scaling. By scaling the output variance inversely with the layer depth, LayerNorm Scaling ensures that all layers, including deeper ones, contribute meaningfully to training. Our experiments show that this simple modification improves performance, reduces resource usage, and stabilizes training across various model sizes. LayerNorm Scaling is easy to implement, hyperparameter-free, and provides a robust solution to enhance the efficiency and effectiveness of LLMs.
8 Impact Statement
This paper introduces the Curse of Depth in LLMs to call attention to the AI community an urgent but often overlooked phenomenon that nearly half of layers in modern LLMs are not as effective as we expect. The impact of this phenomenon is large that a significant amount of resources used to train LLMs are somehow wasted. We further introduce LayerNorm Scaling to ensure that all layers contribute meaningfully to model training. The result is a significant improvement in model efficiency, enabling better performance with fewer computational resources and training tokens. This innovation not only enhances LLM effectiveness across a variety of tasks but also reduces the environmental and financial costs of training large-scale models, making LLM development more sustainable and accessible. LayerNorm Scaling presents a simple, hyperparameter-free solution that can be easily adopted, offering immediate practical benefits to the AI research community.
Appendix A Proofs of the Theorems
A.1 Proof of Lemma 1
@@
Given Equation (2) from [Takase et al., 2023b], we have:
@@
Based on our Assumption 1, let $\mathrm{Var}(\mathrm{Attn}(\mathrm{LN}(x_{\ell})))=\sigma_{\text{Attn}}^{2}$. Then we can write:
@@ A.1.1 Variance of the Attention
$\mathrm{Attn}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$
@@
$\mathrm{Attn}(Q,K,V)=AV.$
To simplify the analysis, we make the following additional assumptions: The softmax output $A$ is approximately uniform, meaning each element of $A$ is roughly $1/n$, where $n$ is the number of keys/values. Given this assumption, the variance of the attention is:
$\mathrm{Var}(\mathrm{Attn}(Q,K,V))\sim\mathrm{Var}(AV)=\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}(V_{i})=\frac{1}{n}\cdot n\cdot\sigma_{V}^{2}=\sigma_{V}^{2}=\sigma_{W}^{2}.$ (22)
@@ A.1.2 Variance of the Feed-Forward Network
Using the result obtained by Wang et al. [2024], we get:
@@
$\sigma^{2}_{x_{\ell+1}}$ $=\sigma^{2}_{x_{\ell}}+\sigma_{W}^{2}+\sigma_{W}^{4}+\rho_{1}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}+\rho_{2}\cdot\sigma_{x^{\prime}_{\ell}}\cdot\sigma_{W}^{2}$ (26)
$=\sigma^{2}_{x_{\ell}}+\sigma_{W}^{2}+\sigma_{W}^{4}+\rho_{1}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}+\rho_{2}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}^{2}+\frac{\rho_{2}\sigma_{W}^{4}}{2\sigma_{x_{\ell}}}+\frac{\rho_{2}^{2}\sigma_{W}^{3}\sigma_{x_{\ell}}}{2}$
@@
Therefore, the increasing lower bound for ${\sigma^{2}_{x_{\ell}}}$ must grows faster than a linear function. So, the increasing of variance is sub-exponential.
A.2 Proof of Theorem 1
@@
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$ (31)
Here, $h$ denotes the number of heads, $s$ is the sequence length, and $d$, $d_{\mathrm{FFN}}$, and $d_{\mathrm{head}}$ are the dimension of the embedding, FFN layer and multi-head attention layer, respectively. The standard deviation of $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{\mathrm{FFN}}$ at layer $\ell$ is $\sigma$ based on Assumption 1.
@@ A.2.1 Proof of Lemma 31
Our derivation follows results in [Takase et al., 2023b], specifically Equation (7), which provides an upper bound on the norm of $\frac{\partial y_{\ell}}{\partial x_{1}}$ as:
@@
Let $\sigma_{w1_{\ell}}$ and $\sigma_{w2_{\ell}}$ be the standard deviations of $W^{1}_{\ell}$ and $W^{2}_{\ell}$, respectively. From Assumption 1, the spectral norms of $W^{1}_{\ell}$ and $W^{2}_{\ell}$ are given by their standard deviations and dimensions [Vershynin, 2018], so wo have:
@@
From [Vershynin, 2018], we know that:
@@
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$ (40)
@@ A.2.2 Analysis of the upper bound
As discussed in [Takase et al., 2023b], $\sigma$ should be sufficiently small, and the standard deviation, $\sigma_{x^{\prime}_{\ell}}$ or $\sigma_{x_{\ell}}$ should satisfy the condition $\sigma^{2}\ll\sigma_{x^{\prime}_{\ell}}$ to maintain the lazy training scheme. Thus, we obtain the following bound for the product over $\ell$ from 1 to $L$:
@@
Equation (31) can be expressed in a traditional product form [Whittaker and Watson, 1996] for $\sigma_{x_{\ell}}$:
@@ A.3 Proof of Lemma 2
Folloing the same analysis as before, we scale the Attention and FFN sub-layers, yielding:
@@
$\log({\sigma^{2}_{x_{\ell}}})$ $=\log\left(\sigma_{x_{1}}^{2}\prod_{k=1}^{\ell-1}\left(1+\sqrt{\frac{1}{{\ell\sigma^{2}_{x_{k}}}}}\right)\right)=\sum_{k=1}^{\ell-1}\log\left(1+\sqrt{\frac{1}{{\ell\sigma^{2}_{x_{k}}}}}\right)+\log(\sigma_{x_{1}}^{2})$ (50)
$\geq\sum_{k=1}^{\ell-1}\Bigl{(}\sqrt{\frac{1}{{\ell\sigma^{2}_{x_{k}}}}}-\frac{1}{2}\left(\sqrt{\frac{1}{{\ell\sigma^{2}_{x_{k}}}}}\right)^{2}\Bigr{)}+\log(\sigma_{x_{1}}^{2}).$
@@
$\sigma^{2}_{x_{\ell}}=\sigma_{x_{1}}^{2}\prod_{k=1}^{\ell-1}\left(1+\frac{1}{\ell^{2-\epsilon/2}}\right)\sim\exp\left(\sum_{k=1}^{\ell-1}\frac{1}{k^{2-\epsilon/2}}\right)\sim\exp\left(\frac{\ell^{\epsilon/2-1}-1}{\epsilon/2-1}\right)\leq\Theta(\ell^{(2-\epsilon)})\leq\Theta(\ell^{2}).$ (52)
Here $\epsilon$ is a small constant with $40<\epsilon\leq 1/$ 4. Therefore, the increasing upper bound of ${\sigma^{2}_{x_{\ell}}}$ is slower than the $\ell^{3}$ function, leading to:
@@ A.4 Proof of Theorem 17
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\ell\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\ell\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$ (55)
Equation (55) is a traditional product form [Whittaker and Watson, 1996] for $\sigma_{x_{\ell}}$. After scaling, it becomes:
@@
Appendix B Training Loss Curve
We report the training loss curve of Pre-LN and LayerNorm Scaling in Figure 5.
Figure 5: Training loss of LLaMA-1B with Pre-LN and LayerNorm Scaling.
Appendix C Variance Growth in Pre-LN Training

### v2

@@ 7 Related Work
Ineffectiveness of Deeper Layers in Transformers. The ineffectiveness of deep layers in LLMs has been previously reported. Yin et al. (2024) found that deeper layers of LLMs can tolerate significantly higher levels of pruning compared to shallower layers, achieving high sparsity. Similarly, Gromov et al. (2024) and Men et al. (2024) demonstrated that removing early layers causes a dramatic decline in model performance, whereas removing deep layers does not. Lad et al. (2024) showed that the middle and deep layers of GPT-2 and Pythia exhibit remarkable robustness to perturbations such as layer swapping and layer dropping. Recently, Li et al. (2024a) highlighted that early layers contain more outliers and are therefore more critical for fine-tuning. While these studies effectively highlight the limitations of deep layers in LLMs, they stop short of identifying the root cause of this issue or proposing viable solutions to address it.
Layer Normalization in Language Models. LN (Ba, 2016) was initially applied after the residual connection in the original Transformer (Vaswani, 2017), which is known as Post-LN. Later on, Pre-LN (Baevski and Auli, 2019; Dai et al., 2019; Nguyen and Salazar, 2019) dominated LLMs, due to its compelling performance and stability (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023; Bi et al., 2024). Prior works have studied the effect of Pre-LN and Post-LN. Xiong et al. (2020) proves that Post-LN tends to have larger gradients near the output layer, which necessitates smaller learning rates to stabilize training, whereas Pre-LN scales down gradients with the depth of the model, working better for deep Transformers. Wang et al. (2019) empirically confirmed that Pre-LN facilitates stacking more layers and Post-LN suffers from gradient vanishing. The idea of connecting multiple layers was proposed in previous works (Bapna et al., 2018; Dou et al., 2018; Wang et al., 2019). Admin introduces additional parameters to control residual dependencies, stabilizing Post-LN. DeepNorm (Wang et al., 2024) enables stacking 1000-layer Transformers by upscaling the residual connection before applying LN. Additionally, Ding et al. (2021) proposed Sandwich LayerNorm, normalizing both the input and output of each transformer sub-layer. Takase et al. (2023a) introduced B2T to bypass all LN except the final one in each layer. Li et al. (2024b) recently combines Post-LN and Pre-LN to enhance the middle layers. Zhu et al. (2025b) introduces Dynamic Tanh (DyT) as a normalization-free alternative in Transformers, delivering comparable performance. Zhuo et al. (2025) proposes HybridNorm, a hybrid normalization scheme combining QKV normalization with Post-Norm FFN to stabilize training in deep transformers. De and Smith (2020) also states that normalized residual blocks in deep networks are close to the identity function and proposes SkipInit to remove normalization by introducing a learnable scalar multiplier on the residual branch initialized to $1/\sqrt{L}$. Our experiments suggest that SkipInit’s learnable parameter does not improve performance and sometimes harms training.
@@ 8 Conclusion
In this paper, we re-introduce the concept of the Curse of Depth in LLMs, highlighting an urgent yet often overlooked phenomenon: nearly half of the deep layers in modern LLMs are less effective than expected. We discover the root cause of this phenomenon is Pre-LN which is widely used in almost all modern LLMs. To tackle this issue, we introduce LayerNorm Scaling. By scaling the output variance inversely with the layer depth, LayerNorm Scaling ensures that all layers, including deeper ones, contribute meaningfully to training. Our experiments show that this simple modification improves performance, reduces resource usage, and stabilizes training across various model sizes. LayerNorm Scaling is easy to implement, hyperparameter-free, and provides a robust solution to enhance the efficiency and effectiveness of LLMs.
Appendix A Proofs of the Theorems of curse of depth
A.1 Proof of Lemma 3.2
@@ Appendix A Proofs of the Theorems of curse of depth
Given Equation (2) from Takase et al. (2023b), we have:
@@
Based on our Assumption 3.1, let $\mathrm{Var}(\mathrm{Attn}(\mathrm{LN}(x_{\ell})))=\sigma_{\text{Attn}}^{2}$. Then we can write:
@@ A.1.1 Variance of the Attention
$\mathrm{Attn}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V.$ (24)
@@
$\mathrm{Attn}(Q,K,V)=AV.$ (25)
Lemma A.1 (Ledoux (2001)).
Let $\{X_{i}\}_{i=1}^{N}$ be independent and identically distributed random variables with mean $m$ and variance $\sigma^{2}<\infty$. Define the softmax weights $p_{i}=\frac{e^{X_{i}}}{\sum_{j=1}^{N}e^{X_{j}}},\quad\text{and let}\quad p=(p_{1},\ldots,p_{N}).$ Then, as $N\to\infty$, with high probability, the softmax vector $p$ concentrates around the uniform distribution on $N$ elements. In particular,
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\sum_{i=1}^{N}\left(p_{i}-\frac{1}{n}\right)^{2}\right]=0,$ (26)
which implies that the softmax output becomes asymptotically indistinguishable, in expectation, from the uniform distribution.
According to the above lemma, to simplify the analysis, we make the following additional assumptions: The softmax output $A$ is approximately uniform, meaning each element of $A$ is roughly $1/n$, where $n$ is the number of keys/values. Given this assumption, the variance of the attention is:
$\mathrm{Var}(\mathrm{Attn}(Q,K,V))\sim\mathrm{Var}(AV)=\frac{1}{n}\sum_{i=1}^{n}d_{\mathrm{head}}\mathrm{Var}(V_{i})=\frac{1}{n}\cdot n\sigma_{V}^{2}\cdot d_{\mathrm{head}}=d_{\mathrm{head}}\sigma_{V}^{2}=\sigma_{W}^{2}d.$ (27)
@@ A.1.2 Variance of the Feed-Forward Network
Using the result obtained by Wang et al. (2024), we get:
@@
$\sigma^{2}_{x_{\ell+1}}$ $=\sigma^{2}_{x_{\ell}}+\sigma_{W}^{2}+\sigma_{W}^{4}d^{2}+\rho_{1}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}+\rho_{2}\cdot\sigma_{x^{\prime}_{\ell}}\cdot\sigma_{W}^{2}d$ (31)
$=\sigma^{2}_{x_{\ell}}+\sigma_{W}^{2}+\sigma_{W}^{4}d^{2}+\rho_{1}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}+\rho_{2}\cdot\sigma_{x_{\ell}}\cdot\sigma_{W}^{2}d+\frac{\rho_{2}\sigma_{W}^{4}d^{2}}{2\sigma_{x_{\ell}}}+\frac{\rho_{2}^{2}\sigma_{W}^{3}d\sigma_{x_{\ell}}}{2}$
@@
∎ Therefore, the increasing lower bound for ${\sigma^{2}_{x_{\ell}}}$ must grow faster than a linear function. So, the increase of variance is sub-exponential. A large increase in such bound will lead to gradient spikes, which can connect to previous studies in Huang et al. (2025a, b).
A.2 Proof of Theorem 3.3
@@
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}$ (36)
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
Here, $h$ denotes the number of heads, $s$ is the sequence length, and $d$, $d_{\mathrm{FFN}}$, and $d_{\mathrm{head}}$ are the dimension of the embedding, FFN layer and multi-head attention layer, respectively. The standard deviation of $W_{Q}$, $W_{K}$, $W_{V}$, and $W_{\mathrm{FFN}}$ at layer $\ell$ is $\sigma$ based on Assumption 3.1.
@@ A.2.1 Proof of Lemma 36
Our derivation follows results in Takase et al. (2023b), specifically Equation (7), which provides an upper bound on the norm of $\frac{\partial y_{\ell}}{\partial x_{1}}$ as:
@@
Let $\sigma_{w1_{\ell}}$ and $\sigma_{w2_{\ell}}$ be the standard deviations of $W^{1}_{\ell}$ and $W^{2}_{\ell}$, respectively. From Assumption 3.1, the spectral norms of $W^{1}_{\ell}$ and $W^{2}_{\ell}$ are given by their standard deviations and dimensions (Vershynin, 2018), so wo have:
@@
From Vershynin (2018), we know that:
@@
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}$ (45)
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
@@ A.2.2 Analysis of the Upper Bound
As discussed in Takase et al. (2023b), $\sigma$ should be sufficiently small, and the standard deviation, $\sigma_{x^{\prime}_{\ell}}$ or $\sigma_{x_{\ell}}$ should satisfy the condition $\sigma^{2}\ll\sigma_{x^{\prime}_{\ell}}$ to maintain the lazy training scheme. Thus, we obtain the following bound for the product over $\ell$ from 1 to $L$:
@@
Equation (36) can be expressed in a traditional product form Whittaker and Watson (1996) for $\sigma_{x_{\ell}}$:
@@
Following the same analysis as before, we scale the Attention and FFN sub-layers, yielding:
@@
$\sigma^{2}_{x_{\ell}}=\sigma_{x_{1}}^{2}\Theta\Big{(}\prod^{\ell-1}_{k=1}\Big{(}1+\frac{1}{\sqrt{k}\sigma_{x_{k}}}\Big{)}\Big{)},$ (55)
We can also obtain a similar result for $\sigma^{2}_{x^{\prime}_{\ell}}$.
@@
$\log({\sigma^{2}_{x_{\ell}}})$ $=\log\left(\sigma_{x_{1}}^{2}\prod_{k=1}^{\ell-1}\left(1+\sqrt{\frac{1}{{k\sigma^{2}_{x_{k}}}}}\right)\right)=\sum_{k=1}^{\ell-1}\log\left(1+\sqrt{\frac{1}{{k\sigma^{2}_{x_{k}}}}}\right)+\log(\sigma_{x_{1}}^{2})$ (56)
$\geq\sum_{k=1}^{\ell-1}\Bigl{(}\sqrt{\frac{1}{{k\sigma^{2}_{x_{k}}}}}-\frac{1}{2}\left(\sqrt{\frac{1}{{k\sigma^{2}_{x_{k}}}}}\right)^{2}\Bigr{)}+\log(\sigma_{x_{1}}^{2}).$
@@
$\sigma^{2}_{x_{\ell}}$ $=\sigma_{x_{1}}^{2}\prod_{k=1}^{\ell-1}\left(1+\frac{1}{k^{2-\epsilon/2}}\right)\sim\exp\left(\sum_{k=1}^{\ell-1}\frac{1}{k^{2-\epsilon/2}}\right)\sim\exp\left(\frac{\ell^{\epsilon/2-1}-1}{\epsilon/2-1}\right)$ (58)
$\leq\Theta(\ell^{(2-\epsilon)})\leq\Theta(\ell^{2}).$
Here $\epsilon$ is a small constant with $11/2\leq\epsilon<$ 1. Therefore, the increasing upper bound of ${\sigma^{2}_{x_{\ell}}}$ is slower than the $\ell^{3}$ function, leading to:
@@ A.4 Proof of Theorem 17
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\Big{(}\Big{(}1+\frac{\sigma^{2}}{\ell\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\Big{)}$ (61)
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\ell\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
Equation (61) is a traditional product form Whittaker and Watson (1996) for $\sigma_{x_{\ell}}$. After scaling, it becomes:
@@
A.5 Proof of theorem 4.3
Proof.
We start with the Equation (3) and the chain‐rule:
$\frac{\partial\mathcal{L}}{\partial W_{1}}=\frac{\partial\mathcal{L}}{\partial y_{L}}\frac{\partial y_{L}}{\partial x_{1}}\frac{\partial x_{1}}{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}\frac{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}{\partial\mathrm{Attn}(x_{1})}\frac{\partial\mathrm{Attn}(x_{1})}{\partial W_{1}}.$ (66)
where $\mathcal{L}$ is the loss function and $\frac{\partial\mathcal{L}}{\partial y_{L}}$ only relates to the composition of loss function. So we only consider:
$\frac{\partial y_{L}}{\partial W_{1}}=\frac{\partial y_{L}}{\partial x_{1}}\frac{\partial x_{1}}{\partial\mathrm{LN}(x_{1})}\frac{\partial\mathrm{LN}(x_{1})}{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}\frac{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}{\partial\mathrm{Attn}(x_{1})}\frac{\partial\mathrm{Attn}(x_{1})}{\partial W_{1}}.$ (67)
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}\leq\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\cdot\left\|\frac{\partial x_{1}}{\partial\mathrm{LN}(x_{1})}\right\|_{2}\cdot\left\|\frac{\partial\mathrm{LN}(x_{1})}{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}\right\|_{2}\cdot\left\|\frac{\partial\mathrm{Attn}(\mathrm{LN}(x_{1}))}{\partial\mathrm{Attn}(x_{1})}\right\|_{2}\cdot\left\|\frac{\partial\mathrm{Attn}(x_{1})}{\partial W_{1}}\right\|_{2}.$ (68)
We know that:
$\left\|\frac{\partial x_{1}}{\partial\operatorname{LN}(x_{1})}\right\|_{2}=\sigma_{x_{1}}.$ (69)
The paper Vershynin (2018) tells us that:
$\left\|\frac{\partial\operatorname{Attn}(\operatorname{LN}(x_{1}))}{\partial\operatorname{LN}(x_{1})}\right\|_{2}=\|W^{O}_{1}J^{Z}_{1}\|_{2}.$ (70)
Now We want to calculate by writing the multi‐head attention (MHA) operation. Although various formulations exist, one common definition of MHA is as follows. For a given input x we compute
$Q=xW^{Q},\quad K=xW^{K},\quad V=xW^{V},$ (71)
and then for each head (we assume head index $i$ and $d_{k}$ the per‐head dimension) $\mathrm{head}_{i}=\operatorname{softmax}\Bigl{(}\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{k}}}\Bigr{)}V_{i}.$ The outputs of all h heads are then concatenated and projected with $W^{O}$: $\operatorname{Attn}(x)=\Bigl{[}\mathrm{head}_{1},\ldots,\mathrm{head}_{h}\Bigr{]}W^{O}.$
$Q=xW_{1},$ and then through $\mathrm{head}=\operatorname{softmax}\Bigl{(}\frac{QK^{\top}}{\sqrt{d_{k}}}\Bigr{)}V,$ so that (ignoring the outer projection $W^{O}$) we have
$\operatorname{Attn}(x)\sim\operatorname{softmax}\Bigl{(}\frac{(xW_{1})K^{\top}}{\sqrt{d_{k}}}\Bigr{)}V.$ (72)
So we have:
$\left\|\frac{\partial\operatorname{Attn}(x)}{\partial W_{1}}\right\|_{2}\leq\frac{\|x\|_{2}}{\sqrt{d_{k}}},$ (73)
That is, writing it out explicitly,
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}\leq\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\sigma_{x_{1}}\|W^{O}_{1}J^{Z}_{1}\|_{2}\frac{\|x\|_{2}}{\sqrt{d}}.$ (74)
We know Equation (44):
Assume that $x\in\mathbb{R}^{n}$ is distributed as a multivariate normal with mean 0 and covariance $\sigma^{2}_{1}$. Then $\sum_{i=1}^{n}y_{i}^{2}$ follows a $\chi^{2}$ distribution with $n$ degrees of freedom. Thus,
$\mathbb{E}\left[\|x\|_{2}\right]=\sqrt{\sigma^{2}_{1}}\mathbb{E}\left[\sqrt{\chi^{2}_{n}}\right]=\sqrt{\sigma^{2}_{1}}\sqrt{2}\frac{\Gamma\Bigl{(}\frac{n+1}{2}\Bigr{)}}{\Gamma\Bigl{(}\frac{n}{2}\Bigr{)}}.$ (76)
For large $n$, the chi-square distribution is concentrated around its mean ($n$) and one often approximates
$\|x_{1}\|_{2}\sim\sqrt{\sigma^{2}_{1}d}.$ (77)
$\|W^{O}\|_{2}\leq\sigma\Bigl{(}\sqrt{d}+\sqrt{hd_{\mathrm{head}}}\Bigr{)}.$ (78)
Combine above together we have:
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}\leq\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}(\sigma_{x_{1}})^{2}\|W^{O}_{1}\|_{2}h\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$ (79)
This is the desired upper bound for $\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}$.
Then we substitute the Equation (36) into the bound: and substituting the bounds for $\|W^{O}_{1}\|_{2}$ and $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ into the original inequality yields
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}$ $\leq\left(\prod_{\ell=1}^{L-1}A_{\ell}\right)(\sigma_{x_{1}})^{2}\sigma\left(\sqrt{d}+\sqrt{hd_{\mathrm{head}}}\right)h$ (80)
×$\quad\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
That is our final upper bound. For clarity, we summarize the answer:
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}\leq$ $\Big{(}\prod_{\ell=1}^{L-1}\left(1+\frac{\sigma^{2}}{\sigma_{x^{\prime}_{\ell}}(\sqrt{d}+\sqrt{d_{\mathrm{FFN}}})^{2}}\right)\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)$ (81)
×$\times\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\left(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{\frac{d_{\mathrm{head}}}{d}}\right)\right)\Big{)}\Big{)}$
×$\quad\times(\sigma_{x_{1}})^{2}\sigma\left(\sqrt{d}+\sqrt{hd_{\mathrm{head}}}\right)h$
×$\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
So we find our that it is in the following form:
$\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}\leq\prod_{l=1}^{L-1}\left(1+\frac{1}{\sigma_{x_{\ell}}}A^{\prime}+\frac{1}{\sigma_{x_{\ell}}^{2}}B^{\prime}\right),$ (82)
Appendix B Variance Growth in Pre-LN Training